OpenStudy (anonymous):

Consider a rope of mass M and length L rotating on top of a table in a horizontal plane. One of the ends of the rope is fixed and does not move. The rope moves in a circle with angular frequency w. What is the kinetic energy of the rope?

OpenStudy (michele_laino):

we can model your problem as below:  now I consider an infinitesimal part of the rope, whose length is \(dx\), it mass is: \[dm = \lambda dx\] where \(\lambda=M/L\) The kinetic energy of this infinitesimal part of rope is: \[dE = \frac{1}{2}dm{v^2} = \frac{1}{2}dm{\omega ^2}{x^2} = \frac{1}{2}\lambda {\omega ^2}{x^2}dx\] being \(v=\omega x\)

OpenStudy (michele_laino):

 the requested kinetic energy, is given by the subsequent integral: \[E = \int_0^L {dE} = \int_0^L {\frac{1}{2}\lambda {\omega ^2}{x^2}dx} \] please complete